if the frequency of an element is 400Hz and a=b=1,then find the element using mosley equation

Grade:

2 Answers

Godfrey Classic Prince
12 years ago

Dear apoorv mishra,

Mosely's Equation is √v = a(Z-b)

Here v is the frequency, a and b are constants !!

Now according to the given information here a=b=1 & v=400 Hz so substituting th values in the equation abouve we get,

√400 = 1(Z-1)

20 = Z-1

Therefore Z = 21 which is Scandium !!

The answer therefore is Scandium .

 

Hope this helped you ..!!

If you loved the answer then please do not forget to approve it by clicking on "Yes" below..!!!

Prashant Goyal
12 years ago

√v=a(z-b)

Therefore z=21 it is scandium